# Flow: A Modular Approach to Automated Agentic Workflow Generation

### v1

@@ Abstract
Multi-agent frameworks powered by large language models (LLMs) have demonstrated great success in automated planning and task execution. However, the effective adjustment of Agentic workflows during execution has not been well-studied. A effective workflow adjustment is crucial, as in many real-world scenarios, the initial plan must adjust to unforeseen challenges and changing conditions in real-time to ensure the efficient execution of complex tasks. In this paper, we define workflows as an activity-on-vertex (AOV) graphs. We continuously refine the workflow by dynamically adjusting task allocations based on historical performance and previous AOV with LLM agents. To further enhance system performance, we emphasize modularity in workflow design based on measuring parallelism and dependence complexity. Our proposed multi-agent framework achieved efficient sub-task concurrent execution, goal achievement, and error tolerance. Empirical results across different practical tasks demonstrate dramatic improvements in the efficiency of multi-agent frameworks through dynamic workflow updating and modularization.
@@ 1 Introduction
Large Language Models (LLMs) [17, 26] show remarkable abilities to understand and generate human-like text. Recent advances have significantly enhanced their capability to emulate human reasoning [18], indicating a promising future for LLM-based reasoning. With the powerful ability to deal with a variety of natural language processing tasks, these models underpin a wide range of applications, from conversational agents [25] and content creation tools [24] to advanced analytics and decision-making systems [15, 20]. Building upon this foundation, a key advancement is the development of multi-agent systems empowered by LLMs [9, 8, 7, 23, 21, 4] where multiple LLM-based agents collaborate to address the same task, leveraging their collective reasoning and planning abilities to automate and optimize task execution processes.
@@
Existing LLMs-based multi-agent systems define LLM as an agent and agents are collaborated with each others via manually designed or LLM-generated prompts. Specifically, MetaGPT [7] focuses on programming tasks by leveraging Standardized Operating Procedures (SOPs) [22, 5, 2]. It predefined distinct roles such as product manager, project manager, and engineer. For each role, an LLM agent is initialized, and these agents follow a strict and sequential workflow to execute sub-tasks. CAMEL [8] is designed to complete a variety of tasks. It requires users to predefine two agents. These agents interact and execute tasks sequentially, with each agent taking on specific responsibilities. AutoGen [23] is also aimed at completing diverse tasks. Unlike CAMEL, AutoGen can automatically create an agent list with different roles based on the task requirements. These agents execute tasks sequentially following the order in the list.
@@
Building upon the strengths of current multi-agent systems, our work aims to further improve existing general-purpose multi-agent systems by enabling dynamically updating workflows during task execution and encouraging modularity in workflows when planning the workflows.
@@
Figure 1: Comparative evaluations among four frameworks—AutoGen, CAMEL, MetaGPT, and Flow (ours)—across two tasks, present notable differences in performance. For the left task, AutoGen, CAMEL, and MetaGPT only managed to produce basic designs lacking in completeness while Flow excelled by creating a fully developed and well-structured website. For the right task, Flow demonstrated superior capability by successfully generating a working game with a clear and intuitive interface while the other frameworks struggled to deliver fully functional code.
@@
Specifically, dynamic updating workflow allows to adjust sub-task allocations and agent roles in real-time based on ongoing performance feedback and changing conditions. This capability ensures that the system remains responsive and efficient even when faced with unexpected obstacles. For instance, if an agent encounters a roadblock in data preprocessing, dynamic updating allows the system to reassign the sub-task to another agent or introduce a new sub-task to overcome the challenge. This adaptability is essential for maintaining robustness and ensuring the seamless execution of complex tasks.
@@
Modularization in system design involves dividing a system into separate, independently operating modules, each responsible for specific functionalities [1]. A highly modularized system allows each module to be developed, managed, and executed in isolation, which simplifies system design and enhances adaptability. In our context, modularization refers to the decomposition of a complex task into smaller, interchangeable sub-task modules. A highly modularized workflow enables sub-tasks to execute concurrently, without bottlenecks from other parts of the workflow. It directly improves the operational efficiency of multi-agent frameworks. In addition, modularity dramatically enhances the ease of dynamic updating. When workflows are highly modularized, the dependency complexity between sub-tasks is small. Therefore, updating one sub-task does not necessitate changes in others, allowing for small adjustments. For instance, if an agent responsible for data preprocessing encounters an unexpected obstacle, the system can dynamically introduce a new sub-task to address the issue with little influence on the rest of the workflow.
@@
In this paper, we have improved existing multi-agent systems by fulfilling modularity and dynamic updating workflow. Our system allows agents to run their sub-tasks in a parallel manner while enabling effective dynamic updates to workflows simultaneously by formulate the entire workflow as an Activity-on-Vertex (AOV) graph, which is a directed acyclic graph (DAG) where each sub-task is represented as a node with its status and generated logs, and the directed edges capture dependencies between sub-tasks. To encourage a modularized workflow design from the beginning, we generate multiple candidate AOV graphs for the task. These candidates are then evaluated based on their degree of parallelism and the complexity of their dependencies. The AOV graph with The highest parallelism and lowest dependency complexity is selected.
@@
During task execution, our system dynamically updates the workflow when a sub-task fails (more detail on Fig. LABEL:fig:flowchart: Running and Tracking status). Updating the system involves modifications to task allocations and agent roles based on ongoing performance data and current workflows. As our AOV-based workflow is encouraged to have high modularity, updating one module does not necessitate changes in others, allowing for localized adjustments during workflow updates (more detail on Fig. LABEL:fig:flowchart: Refining). Similar to the initial workflow generation, multiple AOV graphs are generated and the one with the highest parallelism and lowest dependency complexity is selected during the dynamic updates. This iterative workflow refinement process ensures a good capability of adapting to new challenges and evolving objectives throughout task execution without compromising overall performance.
@@
Our key contributions are as follows: 1) We introduce and encourage modularity in multi-agent workflows, emphasizing the design of workflows with high levels of parallelism and reduced dependency complexities. This modular design enhances efficiency, robustness, and scalability by enabling concurrent task execution and minimizing bottlenecks caused by complex interdependencies. 2) We propose a practical multi-agent framework that supports highly flexible updates to the workflow during runtime. Our method enables updates to the entire workflow based on global information, allowing agents to efficiently adapt to unexpected challenges while maintaining system coherence and consistency. 3)Through comprehensive experiments across multiple datasets, we demonstrate significant improvements in both adaptability and efficiency of our multi-agent system compared to existing approaches. The effectiveness of our method is further validated through a series of experimental evaluations.
@@ 2 Related Work
Recent developments in LLM-driven task decision-making have focused on enhancing the reasoning and planning abilities of agents. Previous approaches like ReAct [24] which iteratively generates thoughts and actions based on current observations until task completion. This framework integrates action-taking with reasoning, allowing agents to perform complex tasks in dynamic environments. Reflexion [16] further improves this by incorporating self-reflection, where the agent evaluates and adjusts its reasoning during execution. ADAPT [13] introduces recursive task decomposition, enabling LLM-based agents to break tasks into smaller subtasks, leading to improved task execution flexibility. However, these approaches often overlook dynamic task reallocation, particularly in multi-agent settings, which is where our work extends the current research.
@@
Multi-agent frameworks have long been employed for task execution in distributed environments, with recent advancements leveraging LLMs to enhance coordination and decision-making. Current frameworks like MetaGPT [7] and CAMEL [8] use structured workflows where multiple agents collaborate to accomplish complex tasks. However, these frameworks often rely on static workflows, which limit their ability to adapt dynamically to changes in the task environment. Recent works like AutoGen [23] address this limitation by introducing more flexible agent collaboration mechanisms. Recent works have explored the use of graphs to represent workflows in multi-agent systems. DyLAN [14] and MACNET [10] utilize static workflows that remain unchanged during execution. GPTSwarm [27] enhances agent interactions but maintains a fixed agent topology, which may limit flexibility in task planning. DataInterpreter [6] updates workflows primarily in response to execution failures in subtasks, adjusting subsequent tasks while leaving completed tasks unchanged. In contrast, our method encourages modularity and facilitates highly flexible modifications to the workflow during runtime, including updates to the agent topology. This capability allows our system to revise and optimize all tasks based on globally generated information, addressing both execution failures and any deficiencies in achieving the overall objectives.
@@ 3 Method
Our proposed framework enhances multi-agent frameworks powered by LLMs by introducing modularity and dynamic workflow updating. This section details how we achieve these features.
@@
Activity on Vertex (AOV) graph is a type of directed acyclic graph (DAG) where vertices represent tasks and edges denote precedence relations [3]. AOV Graphs are crucial in project scheduling and management [11, 19], helping planners visualize dependencies and sequence tasks efficiently.
Inspired by that, we define Multi-Agent workflow as an AOV Graph where vertices represent sub-tasks, with its edges denoting dependencies between these sub-tasks. Let $G=(V,E,A)$ denote the AOV Graph, where $V$ is the set containing all sub-tasks (nodes), $E\subseteq V\times V$ represents the set of directed edges indicating sub-task dependencies, and $A$ represents a set of agents for all sub-tasks. Each agent $a_{j}\in A$ is associated with a role $s_{j}$ and is responsible for executing a subset of tasks $\mathcal{T}_{j}\subseteq V$. We also generate sub-tasks and each directed edge $e_{ij}=(v_{i},v_{j})\in E$ indicates that sub-task $v_{i}$ must be completed before sub-task $v_{j}$ can be started.
Note that AutoGen [23] also automatically generates sub-tasks and agents. However, the sub-tasks are designed to be executed sequentially. For Flow, we allow for the generation of complementary sub-tasks that can run in parallel. This distinction enhances our system’s ability to handle multiple tasks simultaneously, which reduces overall process time and increases efficiency.
@@
Modularity in system design [1] involves dividing a system into separate, independently operating modules, each responsible for specific functionalities, allowing focus on individual components without affecting the entire system. In the context of workflows, we advocate for the creation of sub-tasks that can be executed independently. Modularity is essential for scalability and flexibility in workflows. By reducing dependency complexity, the system can more easily adapt to changes, such as the introduction of new tasks or the reassignment of existing ones, without requiring extensive restructuring. In below theorem, we show that introducing additional dependencies in a workflow reduces the expected success rate of subtasks.
@@
Consider two topologically sorted workflows $A$ and $B$ each consisting of $N$ subtasks according to their execution order. Suppose:
@@
    (Random fail probability) Each subtask $t\in\mathcal{T}$ fails with probability $p_{f}$, where $10<p_{f}<$ 1.
@@
    (Additional dependency in Workflow B) There exists at least one subtask $t^{*}\in\mathcal{T}$ and a subtask $a\in\mathcal{T}$ such that the set of immediate predecessors (dependencies) of $t^{*}$ in Workflow B is $D_{B}(t^{*})=D_{A}(t^{*})\cup\{a\}$, where $D_{A}(t^{*})$ is the set of immediate predecessors of $t^{*}$ in Workflow A. For all other subtasks $t\neq t^{*}$ $D_{A}(t)\subseteq D_{B}(t)$.
Then the expected number of successfully completed subtasks in Workflow A is strictly greater than in Workflow B: $E[S_{A}]>E[S_{B}].$
To encourage modularity in the generated AOV Graph, we define two quantitative measures that evaluate parallelism and dependency complexity, respectively. Parallelism measures the extent to which tasks can be executed concurrently. Let $S_{t}$ represent the set of tasks executed at step $t$. Let $T$ be the total number of steps (the maximum depth of The DAG). Given an AOV Graph $G=(V,E,A)$, the degree of parallelism at a specific step $t$ is defined as the average ratio of the number of tasks executed in that step to the total number of tasks:
$P_{\text{avg}}=\frac{1}{T}\sum_{t=1}^{T}P(t),\text{ where }P(t)=\frac{|S_{t}|}{|V|}.$
While $P_{\text{avg}}$ provides a measure of parallelism, it is insufficient to fully capture the modularity that arises when sub-tasks can be executed independently. Consider two workflows, both containing the same sub-tasks $\{A,B,C,D\}$. For Workflow 1, the task dependencies are defined as: $A\to C,B\to C,A\to D,B\to D,C\to D$. In contrast, Workflow 2 has dependencies: $A\to C,B\to C,C\to D$. Although both workflows exhibit the same level of parallelism, Workflow 2 is structurally simpler in terms of task dependencies, as it contains fewer edges.
To account for this complexity, we measure the dependency structure by analyzing the degree distribution within the task graph. For each task $v_{i}$, the degree $\deg(v_{i})$ reflects the number of direct connections it has in the graph $G$. The average degree $\bar{d}$ is computed as:
$\bar{d}=\frac{1}{|V|}\sum_{v_{i}\in V}\deg(v_{i}),$
where $|V|$ is the number of tasks (vertices) in the graph. The complexity of task dependencies is then quantified by the standard deviation of the degree distribution:
$C_{\text{dependency}}=\left(\frac{1}{|V|}\sum_{v_{i}\in V}(\deg(v_{i})-\bar{d})^{2}\right)^{\frac{1}{2}}.$
This measure reflects the variability in the number of dependencies each task has, providing insight into the overall complexity of the workflow structure.
Task dependencies alone are insufficient to fully capture the modularity that allows sub-tasks to be executed independently. Consider Workflow 3: $A\to B\to C\to D$, which may have a similar dependency complexity to Workflow 2. However, Workflow 2 provides greater modularity and separation of tasks, highlighting the importance of evaluating both dependency complexity and modularity to fully assess and promote effective workflow designs. Both measures are essential for ensuring that tasks can be executed in parallel while maintaining a modular approach.
@@
Our dynamic updates are designed to be flexible, allowing modifications to task allocations including deletion, addition, editing, rerunning, and reassignment of agents. Without a modularity constraint, such flexibility would be difficult to implement. For instance, subtask dependencies can be very complex, and dynamically changing a task could necessitate redoing many existing tasks or incorporating many new tasks. With modularity, efficiency in our dynamic updating process is dramatically enhanced. Intuitively, when workflows are modular, updating one module does not necessitate changes in others.
We leverage Large Language Models (LLMs) as a global inspector to update an AOV Graph based on global information. Specifically, given task requirements $T$, a prompt for update $P_{\text{update}}$, the current AOV Graph $G^{t}$, and generated data $D^{t}$ containing the status of subtasks and the output of agents for running subtasks, the LLM continuously monitors task progress and dynamically modifies the graph when necessary. Similar to the initialization process, we generate $K$ candidate graphs: $\{G_{1}^{t+1},G_{2}^{t+1},\dots,G_{K}^{t+1}\}=f(P_{\text{update}},T,D^{t})$. We follow the same selection strategy as in initialization which prioritizes the workflow with the highest parallelism score. If after the selection, the graph is not unique, we further select the one with the lowest dependency complexity.
Note that with sufficient data and computational resources, we could further enhance our framework by fine-tuning LLMs with reinforcement learning (RL) for workflow generation. For example, the LLM would be trained to maximize a reward function designed around key performance indicators such as task completion speed, resource utilization, and minimization of workflow disruptions.
@@
Our framework employs a dictionary-based structure, $\tilde{G}$, to efficiently manage and dynamically update workflows within a multi-agent system. This approach represents each task $v$ in the workflow as a key in $\tilde{G}$, with the value being another dictionary that encapsulates various attributes of the task. The structure is specifically defined as:
@@
Each task’s dictionary includes attributes such as the sub-task requirement, current status (e.g., ”not started”, ”in progress”, ”completed”), data relevant to the task, a count of uncompleted parent tasks to manage dependencies, a list of child tasks that depend on the current task’s completion, and the agent assigned to the task. The choice of a dictionary-based structure for our workflow system is driven by its inherent simplicity and flexibility. This structure can be converted directly to JSON, and the organized information is easily readable and summarizable by large language models (LLMs).
Each task’s execution readiness is determined by the attribute ”num_parents_not_completed”. Tasks with a count of zero are eligible to run concurrently, leveraging our system’s capability to handle parallel task execution effectively. Upon the completion of any task, we perform a systematic review to determine if the workflow requires refinement, ensuring that all dependencies are accurately accounted for and that the workflow remains aligned with project goals. Additionally, we do not rely solely on the status and ”num_parents_not_completed” counts reported by agents. These are always double-checked to prevent errors that could arise from inaccurate reporting by agents or unforeseen system anomalies. This rigorous verification process enhances the reliability and integrity of our workflow management system.

### v2

# Flow: Modularized Agentic Workflow Automation

@@ Abstract
Multi-agent frameworks powered by large language models (LLMs) have demonstrated great success in automated planning and task execution. However, the effective adjustment of agentic workflows during execution has not been well studied. An effective workflow adjustment is crucial in real-world scenarios, as the initial plan must adjust to unforeseen challenges and changing conditions in real time to ensure the efficient execution of complex tasks. In this paper, we define workflows as an activity-on-vertex (AOV) graph, which allows continuous workflow refinement by LLM agents through dynamic subtask allocation adjustment based on historical performance and previous AOVs. To further enhance framework performance, we emphasize modularity in workflow design based on evaluating parallelism and dependency complexity. With this design, our proposed multi-agent framework achieves efficient concurrent execution of subtasks, effective goal achievement, and enhanced error tolerance. Empirical results across various practical tasks demonstrate significant improvements in the efficiency of multi-agent frameworks through dynamic workflow refinement and modularization. The code is available at: https://github.com/tmllab/2025_ICLR_FLOW.
@@ 1 Introduction
Large Language Models (LLMs) [19, 31] show remarkable ability to understand and generate human-like text. Recent advances have significantly enhanced their capability to emulate human reasoning [21], indicating a promising future for LLM-based reasoning. With the powerful ability to handle a variety of natural language processing tasks, these models underpin a wide range of applications, from conversational agents [28] and content creation tools [27] to advanced analytics and decision-making systems [17, 23]. Building upon this foundation, a key advancement is the development of multi-agent frameworks empowered by LLM [11, 10, 8, 26, 24, 5, 12] where multiple LLM-based agents collaborate to address complex tasks, leveraging their collective reasoning and planning abilities to automate and optimize task execution processes.
@@
Existing LLM-based multi-agent frameworks define LLM as an agent, and agents collaborate with each other via manually designed or LLM-generated prompts. Specifically, MetaGPT [8] focuses on programming tasks by leveraging Standardized Operating Procedures (SOPs) [25, 6, 3]. It predefined distinct roles such as product manager, project manager, and engineer. For each role, an LLM agent is initialized, and these agents operate within a strict and sequential workflow to execute subtasks. CAMEL [10] can complete a variety of task types. It requires users to pre-define two agents. These agents interact and execute tasks sequentially, each agent taking on specific responsibilities. AutoGen [26] is also aimed at completing diverse tasks. Unlike CAMEL, AutoGen can automatically create an agent list with different roles based on subtask requirements. These agents execute subtasks sequentially following the order in the list.
@@
Building upon the strengths of current multi-agent frameworks, our work aims to further improve existing general-purpose multi-agent frameworks by enabling dynamically updating workflows during task execution and encouraging modularity in workflows when planning the workflows.
@@
Figure 1: Comparative evaluations among four frameworks—AutoGen, CAMEL, MetaGPT, and Flow (ours)—across two tasks, present notable differences in performance. For the left task, AutoGen, CAMEL, and MetaGPT only managed to produce basic designs lacking in completeness while Flow excelled by creating a fully developed and well-structured website. For the right task, Flow demonstrated superior capability by successfully generating a working game with a clear and intuitive interface, while the other frameworks struggled to deliver fully functional code.
@@
Specifically, dynamic updating workflows allow agents to adjust subtask allocations and agent roles in real-time based on ongoing performance feedback and changing conditions. This capability ensures that the system remains responsive and efficient even when faced with unexpected obstacles. For instance, if an agent encounters a roadblock in data preprocessing, the system can reassign this subtask to another agent or introduce a new subtask to resolve the issue. Such adaptability is essential for maintaining robustness and ensuring the seamless execution of complex tasks.
@@
Modularity In system design, involves dividing a system into separate, independently operating modules, each responsible for specific functionalities [2]. In our context, modularity refers to the decomposition of a complex task into smaller, interchangeable subtask modules. A highly modularized workflow enables subtasks to execute concurrently, without bottlenecks from other parts of the workflow and thereby directly improves the operational efficiency of multi-agent frameworks. Furthermore, modularity enhances the ease of dynamic updating. When workflows are highly modularized, the dependency complexity between subtasks is minimal. Therefore, updating one subtask does not affect others, allowing for small workflow adjustments. For example, if an agent responsible for data preprocessing encounters an unexpected obstacle, a system of high modularity can adapt by introducing only one subtask with minimal impact on the rest of the workflow.
@@
In this paper, we enhance existing multi-agent frameworks by achieving modularity and enabling dynamic workflow updates. Our framework allows agents to execute their subtasks in parallel while facilitating efficient workflow updates. This is accomplished by formulating the entire workflow as an Activity-on-Vertex (AOV) graph, which is a directed acyclic graph (DAG) where each subtask is represented as a node with its status and generated logs, while the directed edges capture dependencies between subtasks. To encourage a modular workflow design from the beginning, we generate multiple candidate AOV graphs for the task. These candidates are then evaluated based on their degree of parallelism and the complexity of their dependencies. The AOV graph with the highest parallelism and lowest dependency complexity is selected.
@@
During task execution, our framework continuously checks and refines the workflow, updating it when a subtask fails (see Fig. LABEL:fig:flowchart: Check & Refine). The framework updates subtask allocations and agent roles based on ongoing performance data and current workflow. As our AOV-based workflow encourages high modularity, updating one module does not necessarily affect others, allowing for localized adjustments during workflow updates (see Fig. LABEL:fig:flowchart: Update). Similar to the initial workflow generation, multiple AOV graphs are generated and the one with highest parallelism and lowest dependency complexity is selected during dynamic updates. This iterative workflow refinement process enhances adaptability to new challenges and evolving objectives throughout task execution, ensuring dynamic workflow updates without compromising overall performance.
@@
Our key contributions are as follows: 1) We introduce and encourage modularity in multi-agent workflows, emphasizing the design of workflows with high parallelism and low dependency complexity. This modular design enhances efficiency, robustness, and scalability by enabling concurrent subtask execution and minimizing bottlenecks caused by complex interdependence. 2) We propose a practical multi-agent framework that supports highly flexible updates to the workflow during runtime. Our method enables local updates to the entire workflow based on global information, allowing agents to efficiently adapt to unexpected challenges while maintaining system coherence and consistency. 3)Through comprehensive experiments, we demonstrate significant improvements in both the adaptability and efficiency of our multi-agent framework compared to existing approaches.
@@ 2 Related Work
Recent developments in LLM-based task decision making have focused on improving the reasoning and planning abilities of agents [27, 20, 30, 26, 15, 18, 1]. Previous approaches like ReAct [27] iteratively generate thoughts and actions based on current observations until task completion. This framework integrates action-taking with reasoning, allowing agents to perform complex tasks in dynamic environments. Reflexion [18] further improves this by incorporating self-reflection, where the agent evaluates and adjusts its reasoning during execution. ADAPT [15] introduces recursive task decomposition, enabling LLM-based agents to break tasks into smaller subtasks, which leads to improved task execution flexibility. However, these approaches often overlook dynamic task reallocation, particularly in multi-agent settings, which is where our work extends the current research.
@@
Multi-agent frameworks have long been employed for task execution in distributed environments, with recent advances leveraging LLM to enhance coordination and decision-making [8, 10, 26, 9]. However, existing frameworks often rely on static workflows with limited adaptability to changes in the task environment. DyLAN [12] and MACNET [16] utilize static graphs to represent workflows in multi-agent frameworks; GPTSwarm [32] enhances agent interactions but maintains a fixed agent topology; DataInterpreter [7] updates workflows primarily in response to execution failures in subtasks, adjusting subsequent tasks while leaving completed tasks unchanged; AFlow [29] introduces a dynamic workflow generation framework based on Monte Carlo Tree Search, enabling adaptive adjustments through iterative code modification. This highlights the need for dynamic workflow updates.
@@ 3 Method
Our proposed Flow enhances multi-agent frameworks powered by LLM by introducing modularity and dynamic workflow updating. As depicted in Fig. LABEL:fig:flowchart, given the task requirement, Flow first formulates the initial workflow for execution plan generation and agent allocation. During execution, the workflow is continuously refined and dynamically updated until the task is completed. To maximize system simplicity and flexibility, we design a dictionary-based structure for implementation. In the following, we detail how to achieve these features.
@@
Activity on Vertex (AOV) graph is a type of directed acyclic graph where vertices represent subtasks and edges denote precedence relations [4]. AOV graphs are widely used in project scheduling and management [13, 22], helping planners visualize dependencies and sequence subtasks efficiently.
Inspired by that, we define the multi-agent workflow as an AOV graph where vertices represent subtasks, while edges denote dependencies between subtasks. Let $G=(V,E,A)$ denote the AOV graph, with $V$ the set of all subtasks (vertices), $E\subseteq V\times V$ the set of directed edges indicating subtask dependencies. For example, $e_{ij}=(v_{i},v_{j})\in E$ indicates that the subtask $v_{i}$ must be completed before the subtask $v_{j}$ starts. $A$ represents a set of agents for all subtasks. Each agent $a_{j}\in A$ is associated with a role that is responsible for executing a subset of subtasks $\mathcal{T}_{j}\subseteq V$.
Note that AutoGen [26] also automatically generates subtasks and agents. However, the subtasks are designed to be executed sequentially. For Flow, we allow for the generation of complementary subtasks that can run in parallel. This distinction enhances our framework’s ability to handle multiple subtasks simultaneously, which reduces overall process time and increases efficiency.
@@
Modularity in system design [2] involves dividing a system into separate, independently operating modules, each responsible for specific functionalities, allowing focus on individual components without affecting the entire system. It is essential for scalability and flexibility in workflows. By reducing dependency complexity, the system can more easily adapt to changes, such as the introduction of new tasks or the reassignment of existing ones, without requiring extensive restructuring. Theorem 3.1 demonstrates additional dependencies in a workflow reduce the expected success rate of subtasks. Following this conclusion, Flow  advocates for the creation of subtasks that can be executed independently.
@@
Consider two topologically sorted workflows $A$ and $B$ each consisting of $N$ subtasks according to their execution order. Suppose
@@
    (Random fail probability) Each subtask $v\in\mathcal{T}$ fails with probability $p_{f}$, where $10<p_{f}<$ 1.
@@
    (Additional dependency in Workflow B) There exist at least one subtask $v^{*}\in\mathcal{T}$ and a subtask $b\in\mathcal{T}$ such that the set of immediate predecessors (dependencies) of $v^{*}$ in Workflow B is $D_{B}(v^{*})=D_{A}(v^{*})\cup\{b\}$, where $D_{A}(v^{*})$ is the set of immediate predecessors of $v^{*}$ in Workflow A. For all other subtasks $v\neq v^{*}$ $D_{A}(v)\subseteq D_{B}(v)$.
The expected number of completed subtasks in Workflow A is strictly greater than in Workflow B: $E[S_{A}]>E[S_{B}].$
To encourage modularity in the generated AOV graph, we define two quantitative measures that evaluate parallelism and dependency complexity respectively. Parallelism measures the extent to which subtasks can be executed concurrently. Let $S_{t}$ represent the set of subtasks executed in the $t$ step. Let $T$ be the total number of steps (the maximum depth of the DAG). Given an AOV graph $G=(V,E,A)$, the degree of parallelism overall is defined as the average subtask ratio over steps:
$P_{\text{avg}}=\frac{1}{T}\sum_{t=1}^{T}S_{t}.$
Although $P_{\text{avg}}$ provides a measure of parallelism, it is insufficient to fully capture the modularity that arises when subtasks can be executed independently. Consider two workflows, both containing the same subtasks $\{A,B,C,D\}$. For Workflow 1, the task dependencies are defined as: $A\to C,B\to C,A\to D,B\to D,C\to D$. In contrast, Workflow 2 has dependencies: $A\to C,B\to C,C\to D$. Although both workflows exhibit the same level of parallelism, Workflow 2 is structurally simpler in terms of task dependencies, as it contains fewer edges.
To account for this complexity, we measure the dependency structure by analyzing the degree distribution within the subtask graph. For each subtask $v_{i}$, we define $\deg(v_{i})$ as the number of direct connections it has on the graph $G$. The dependency complexity is quantified by the standard deviation of the number of direct connections:
$C_{\text{dependency}}=\sigma_{\deg(v_{i})}=\sqrt{\frac{1}{|V|}\sum_{v_{i}\in V}(\deg(v_{i})-\bar{d})^{2}}.$
This measure reflects the variability in the number of dependencies each subtask has, providing insight into the overall complexity of the workflow structure.
Task dependencies alone are insufficient to fully capture the modularity that allows subtasks to be executed independently. Consider Workflow 3: $A\to B\to C\to D$, which may have a similar dependency complexity to Workflow 2. However, Workflow 2 provides greater modularity and separation of subtasks, highlighting the importance of evaluating both dependency complexity and modularity to fully assess and promote effective workflow designs. Both measures are essential to ensure that subtasks can be executed in parallel while maintaining a modular approach.
@@
We leverage LLM as a global inspector to continuously monitor task progress and dynamically modify the AOV graph based on global information when necessary. Specifically, given the task requirements prompt $\mathcal{P}$ and the update prompt $\mathcal{P}_{\text{update}}$, the current AOV graph $G^{t}$, and the generated data $D^{t}$ containing the status of subtasks and the output of agents to run subtasks. Similarly to the initialization process, we generate $K$ candidate graphs: $\{G_{1}^{t+1},G_{2}^{t+1},\dots,G_{K}^{t+1}\}=f(\mathcal{P}_{\text{update}},\mathcal{P},D^{t})$. We follow the same selection strategy as in initialization, which prioritizes the workflow with the highest parallelism score and further selects the one with the lowest dependency complexity if multiple graphs share the highest parallelism score.
With the modularity constraint introduced in previous sessions, our dynamic updates can largely fulfill flexibility, allowing modifications to subtask allocations including deletion, addition, editing, rerunning, and reassignment of agents without necessarily affecting other agents or their assigned subtasks. This unique advantage is particularly beneficial when subtask requirements become more challenging, as subtask dependencies can be highly complex.
Note that with sufficient data and computational resources, we could further enhance our framework by fine-tuning LLM with reinforcement learning for workflow generation. For example, the LLM would be trained to maximize a reward function designed around key performance indicators such as task completion speed, resource utilization, and minimization of workflow disruptions.
@@
Our framework employs a dictionary-based structure, $\tilde{G}$, to efficiently manage and dynamically update workflows within a multi-agent framework. Each subtask $v$ in the workflow is represented as a key in $\tilde{G}$, the value being another dictionary that encapsulates various attributes of the subtask. The structure is specifically defined as:
@@
In each $\tilde{G}[v]$, the values of each key are as follows:
  • ”subtask requirement”: the text of the task requirement;
  • ”status”: the current task implementation status e.g. ”not started”, ”in progress”, ”completed”;
  • ”data”: data relevant to this task;
  • ”num_parents_not_completed”: the count of uncompleted parent tasks to manage dependencies;
  • ”child”: a list of child tasks that depend on the current task’s completion;
  • ”agent”: the agent assigned to the task.
This dictionary-based structure can be converted directly to JSON, and the organized information is easily readable and summarizable by LLM, granting our system inherent simplicity and flexibility. Each subtask execution readiness is determined by the attribute ”num_parents_not_completed”. Subtasks with a count of zero are eligible to run concurrently, leveraging our system’s capability to handle parallel subtask execution effectively. Upon completion of each subtask, we perform a systematic review to determine if the workflow requires refinement, ensuring that all dependencies are accurately accounted for and that the workflow remains aligned with project goals. In addition to monitoring the subtask completion by the ”status” and ”num_parents_not_completed” counts reported by agents. Flow also double-checks the completion of each subtask by asking if all the requirements of this subtask are fulfilled. This will largely prevent errors from inaccurate reporting by agents or unforeseen system anomalies. This rigorous verification process enhances the reliability and integrity of our workflow management system.